# Flex3D: Feed-Forward 3D Generation With Flexible Reconstruction Model And Input View Curation

### v2

@@ Abstract
Generating high-quality 3D content from text, single images, or sparse view images remains a challenging task with broad applications. Existing methods typically employ multi-view diffusion models to synthesize multi-view images, followed by a feed-forward process for 3D reconstruction. However, these approaches are often constrained by a small and fixed number of input views, limiting their ability to capture diverse viewpoints and, even worse, leading to suboptimal generation results if the synthesized views are of poor quality. To address these limitations, we propose Flex3D, a novel two-stage framework capable of leveraging an arbitrary number of high-quality input views. The first stage consists of a candidate view generation and curation pipeline. We employ a fine-tuned multi-view image diffusion model and a video diffusion model to generate a pool of candidate views, enabling a rich representation of the target 3D object. Subsequently, a view selection pipeline filters these views based on quality and consistency, ensuring that only the high-quality and reliable views are used for reconstruction. In the second stage, the curated views are fed into a Flexible Reconstruction Model (FlexRM), built upon a transformer architecture that can effectively process an arbitrary number of inputs. FlexRM directly outputs 3D Gaussian points leveraging a tri-plane representation, enabling efficient and detailed 3D generation. Through extensive exploration of design and training strategies, we optimize FlexRM to achieve superior performance in both reconstruction and generation tasks. Our results demonstrate that Flex3D achieves state-of-the-art performance, with a user study winning rate of over 92% in 3D generation tasks when compared to several of the latest feed-forward 3D generative models.
See project page for more immersive 3D results.
@@ 1 Introduction
Fast generation of high-quality 3D contents is becoming increasingly important for video games development (Hao et al., 2021; Sun et al., 2023), augmented, virtual, and mixed reality (Li et al., 2022), robotics (Nasiriany et al., 2024) and many other applications. Recent advances in computer vision and graphics (Mildenhall et al., 2021; Kerbl et al., 2023) and deep learning (Dosovitskiy, 2020; Caron et al., 2021; Oquab et al., 2023), combined with the availability of large datasets of 3D objects (Deitke et al., 2023; 2024; Yu et al., 2023; Chang et al., 2015), have made it possible to learn neural networks that can generate 3D objects from text, single images or a sparse set of views, and to do so in an efficient, feed-forward manner.
A particularly successful family of 3D generators are the ones based on sparse-views reconstruction (Xu et al., 2024c; Siddiqui et al., 2024; Li et al., 2024b; Zhang et al., 2024c; Tang et al., 2024a; Xu et al., 2024b; Xie et al., 2024; Wang et al., 2024c). Compared to single-image reconstructors, multi-view reconstruction models generally produce better 3D assets. This advantage arises because the multi-view images implicitly capture the object geometry much better, substantially simplifying the reconstruction problem. However, to generate a 3D object from text or a single image, one must first synthesize several views of the objects, for example by means of a multi-view diffusion model. These multi-view diffusion models often generate inaccurate and inconsistent views, which are difficult for the reconstruction network to reconcile, and can thus affect the overall quality of the final 3D output (Tang et al., 2024c).
@@
We benchmark our method against state-of-the-art feed-forward models in 3D generation and 3D reconstruction tasks, evaluating performance through a user study and various automated metrics. We achieve the best results in reconstruction tasks across all settings (single-view, four-view, and more-view), as well as in generation tasks. We also conduct a thorough ablation study to assess the impact of our design choices.
In summary, this paper makes the following key contributions: (1) We introduce a new generation pipeline that starts by generating a pool of 2D views of an object and then selects the best ones for 3D reconstruction. (2) We propose FlexRM, a 3D reconstruction network that can efficiently process an arbitrary number of input views with varying viewpoints, obtaining a complete reconstruction of the 3D object which can then be converted to a mixture of 3D Gaussians. (3) We further introduce a new training strategy to enhance the performance and robustness of the 3D reconstructor by simulating imperfections in the input views during training, leveraging the model’s own outputs and injecting noise into the Gaussians.
@@ 2.1 Multi-view Generation
Zero123 (Liu et al., 2023a) first proposed using multi-view data to fine-tune an image diffusion model for generating novel views from a single view, conditioned on camera parameters. Following this approach, subsequent works (Li et al., 2024b; Shi et al., 2023b; Tang et al., 2023; Liu et al., 2023b; Long et al., 2024; Wang & Shi, 2023; Woo et al., 2024; Yang et al., 2024b; Ye et al., 2024; Zhao et al., 2024; Zheng & Vedaldi, 2024; Tang et al., 2024b) largely focused on generating multiple views simultaneously to ensure 3D consistency.
@@ 2.2 Feed-forward 3D Reconstruction and Generation
Recent advances in 3D reconstruction and generation have focused on training feed-forward models that directly output 3D representations without requiring further optimization (Yu et al., 2021; Erkoç et al., 2023; Szymanowicz et al., 2023b; Ren et al., 2023; Lorraine et al., 2023; Xu et al., 2024a; Tochilkin et al., 2024; Zhang et al., 2024a; Jiang et al., 2024; Han et al., 2024a). These feed-forward models offer significant advantages in both reconstruction quality and inference speed.
@@
A representative series of work is LRM (Hong et al., 2024), which learns to generate a tri-plane NeRF (Chan et al., 2022) representation using a transformer network. This approach can receive multiple types of inputs, including single images, text (Xu et al., 2024d), posed sparse-view images (Li et al., 2024b), and unposed sparse-view images (Wang et al., 2024a). Further works (Wei et al., 2024; Siddiqui et al., 2024; Xu et al., 2024b; Wang et al., 2024c; Boss et al., 2024) focused on improving the geometric quality of generated 3D assets. Some (Zou et al., 2023) proposed to combine the tri-plane representation with 3D Gaussian Splatting for more efficient rendering. They suggest using an additional point cloud network to determine the 3D Gaussian position to overcome the tendency of 3D Gaussian to get stuck in local optima.
@@
Another representative series of work (Tang et al., 2024a; Xu et al., 2024c; Zhang et al., 2024c) generates 3D Gaussian points directly through per-pixel aligned Plücker ray embedding and predicts the depth for each pixel (Szymanowicz et al., 2023a), which can then be converted to 3D Gaussian locations. However, such approaches requires the input views to cover a large visible range of the 3D object. Building an intermediate 3D feature representation to regress 3D Gaussian points is also possible (Chen et al., 2024; Zhang et al., 2024b), but these methods still require sparse-view images as input and typically prefer a fixed number of views with fixed viewing angles. However, in 3D generation tasks where sparse input views are generated through a multi-view diffusion model and are not always of high quality, such sparse-view reconstructors tend to produce suboptimal results.
@@
This paper introduces a flexible 3D reconstruction model that combines the strengths of the approaches above. Our tri-plane-based model efficiently generates high-quality 3D Gaussian points directly, without needing additional modules. It also accommodates a variable number of input views, enabling integration with our view selection pipeline for high-quality 3D generation.
@@ 3.1 Candidate View Generation and Selection
Our image/text-to-multi-view-images generator module generates four views of the 3D object from four elevation degrees (-18°, 6°, 18°, and 30°) We utilize the Emu model (Dai et al., 2023), which is pre-trained on a massive dataset of billions of text-image pairs, as our base model. Following prior works (Shi et al., 2023b; Li et al., 2024b; Siddiqui et al., 2024), we fine-tune this model on approximately 130,000 rendered multi-view images. This fine-tuning process enables the model to predict a 2$\times$2 grid of four consistent images, each corresponding to one of the specified elevation angles.
@@
As the two multi-view diffusion models are focused on different aspects (elevation and azimuth) of novel view generation, there are minimal conflicts in the generated views between them. Even so, and despite efforts to improve the quality of the outputs, not all generated views are entirely consistent. Certain views, particularly those from challenging angles like the back, may exhibit suboptimal generation quality, and there can be inconsistencies between different generated views. Including such flawed views as input for 3D reconstruction can significantly degrade the quality of the final 3D asset. Therefore, we introduce a mechanism to filtering poor-quality views. This is done via a novel view selection pipeline which consists of two steps:
(1) Back View Quality Assessment: We employ a multi-view video quality classifier trained to assess the overall quality of generated videos, with particular emphasis on the quality of the back view. This classifier utilized DINO (Oquab et al., 2023) to extract features from the front view and back view of the multi-view video, and subsequently trained a Support Vector Machine (SVM) to classify video quality based on the combined DINO features. The training data consisted of 2000 manually labeled “good” and “bad” Emu-generated video samples.
We apply the quality classifier to the multi-view video to determine whether the back view exhibits reasonable generation quality.
@@
The whole process of view selection can be done in less than a second with a single A100 GPU.
@@ 3.2 Flexible Reconstruction Model (FlexRM)
The position of the 3D Gaussian is expressed as $\mathbf{p}=\alpha\mathbf{p}_{0}+(1-\alpha)\delta\mathbf{p}=\alpha\mathbf{p}_{0}+(1-\alpha)f(\mathbf{p}_{0})$ where $\delta\mathbf{p}$ is a a positional offset output by the MLP and $\alpha=3/4$. These positional offsets $\delta\mathbf{p}$ are constrained to the range of $[-1,1]$ through the application of the tanh activation function. This approach, akin to residual learning, facilitates the optimization process. The multipliers ensure that $\mathbf{p}$ remains within the same range as $\mathbf{p}_{0}$, which prevents Gaussian points from moving beyond the visible boundaries, which would result in their projection falling outside the 2D image plane and consequently providing no gradients for optimization. Furthermore, since $\alpha<1$, this expression biases Gaussians to shift towards the center of the tri-plane grid, where the object is usually located.
@@
Our training dataset comprises multi-view dense renders from an internal dataset analogous to Objaverse. Specifically, for each object, we render 512$\times$512 resolution images from 256 viewpoints, uniformly distributed across 16 azimuth and 16 elevation angles. This process yields approximately 700,000 rendered objects, with 140,000 classified as high-quality. Furthermore, we leverage the Emu-video synthetic dataset (Han et al., 2024a), which consists of 2.7 million synthetic multi-view videos. Each video comprises 16 frames capturing a 360-degree azimuth at a fixed elevation angle.
@@
More details on implementation and training of FlexRM are presented in the Appendix A.
Figure 3: Imperfect Input View Simulation Results. We simulate different kinds of imperfect input views by feeding FlexRM’s output back as input and manipulating 3D Gaussian parameters.
@@ 3.3 Imperfect input view simulation
While FlexRM’s outputs naturally contain small imperfections, we also introduce random perturbations to FlexRM’s generated 3D Gaussian points during step (1) to simulate a wider range of imperfect inputs and promote greater diversity. We sample a randomly sized small cube from the large tri-plane cube and add noise with varying intensities to the 3D Gaussian parameters, excluding rotation. For example, adding noise to positions can simulate part-level 3D inconsistencies, while adding noise to color parameters can simulate color distortions. Adding noise to opacity results in a speckled or streaky appearance, while adding noise to scale leads to a blurring effect. Figure 3 illustrates these effects. During training, each effect is randomly used with a probability of 0.2. This combines them for greater diversity. Please check Appendix A for more details.
@@ 4 Experiments
We leverage 404 deduplicated prompts from DreamFusion (Poole et al., 2022) to conduct an experiment on text-to-3D or single-image-to-3D generation. We compare Flex3D to OpenLRM (He & Wang, 2023; Hong et al., 2024), VFusion3D (Han et al., 2024a), LGM (Tang et al., 2024a), InstantMesh (Xu et al., 2024b), and GRM (Xu et al., 2024c). For GRM, we utilize its provided Instant3D’s multi-view diffusion model to generate input multi-view images, and for InstantMesh, we employ the default Zero123++ (Shi et al., 2023a) for text-to-input multi-view image conversion.
@@
To further evaluate the overall quality of the generated content, we conducted a user study. Participants were presented with pairs of 360° rendered videos—one generated by Flex3D and one by a baseline model—and asked to select their preferred video. We randomly selected 40 prompts for evaluation. The corresponding 40 pairs of generated videos were then independently evaluated by five users, with each user assessing all 40 pairs. For each pair, we collect five results, and the majority preference was recorded as a win rate for Flex3D.
[…]
OpenLRM 1 15.83 0.821 0.209 0.602
VFusion3D 1 19.10 0.827 0.158 0.759
FlexRM 1 21.21 0.862 0.125 0.832
InstantMesh 4 21.33 0.859 0.133 0.809
GRM 4 25.03 0.899 0.102 0.869
FlexRM 4 25.55 0.894 0.074 0.893
FlexRM 8 26.33 0.897 0.069 0.906
FlexRM 16 26.51 0.902 0.068 0.911
Table 2: Reconstruction Performance on the GSO Dataset. FlexRM consistently outperforms other baselines, where it achieves the best results across different input view settings. Increasing the number of input views for FlexRM generally leads to improved reconstruction quality.
Results are also shown in table 1, where a significant number of votes goes to our method for its high quality, regardless of the baselines used for comparison. This shows that our method clearly generates better 3D assets.
@@
In table 2 we report the performance of FlexRM on the GSO reconstruction task, using varying numbers of input views, namely 1, 4, 8, and 16. We compare our results to several baseline methods, including single-view reconstruction models (LRM (Hong et al., 2024; He & Wang, 2023), VFusion3D (Han et al., 2024a)) and sparse-view reconstruction models (InstantMesh (Xu et al., 2024b), GRM (Xu et al., 2024c)). For the single-view setting, we use the input view at 0° azimuth and 10° elevation as input. For the 4-view setting, we use views at 0°, 90°, 270°, and 360° azimuth degrees, all at 10° elevation. For other numbers of views, we heuristically select more views as input. The remaining views are used to compute the reported novel-view synthesis quality.
Overall, FlexRM significantly outperforms baselines in both 1-view and 4-view settings. This improvement is particularly evident in the LPIPS score, a key metric reflecting perceptual quality, which demonstrates substantial gains. Beyond fixed input views, FlexRM is also capable of handling an arbitrary number of input views. We present results for 8-view and 16-view results, both exhibiting progressively stronger performance. Qualitative results are shown in the Appendix D.

### v3

@@ Abstract
Generating high-quality 3D content from text, single images, or sparse view images remains a challenging task with broad applications. Existing methods typically employ multi-view diffusion models to synthesize multi-view images, followed by a feed-forward process for 3D reconstruction. However, these approaches are often constrained by a small and fixed number of input views, limiting their ability to capture diverse viewpoints and, even worse, leading to suboptimal generation results if the synthesized views are of poor quality. To address these limitations, we propose Flex3D, a novel two-stage framework capable of leveraging an arbitrary number of high-quality input views. The first stage consists of a candidate view generation and curation pipeline. In the second stage, the curated views are fed into a Flexible Reconstruction Model (FlexRM), built upon a transformer architecture that can effectively process an arbitrary number of inputs. Through extensive exploration of design and training strategies, we optimize FlexRM to achieve superior performance in both reconstruction and generation tasks. Our results demonstrate that Flex3D achieves state-of-the-art performance, with a user study winning rate of over 92% in 3D generation tasks when compared to several of the latest feed-forward 3D generative models. See project page for more immersive 3D results.
@@ 1 Introduction
Fast generation of high-quality 3D contents is becoming increasingly important for video games development (Hao et al., 2021; Sun et al., 2023), augmented, virtual, and mixed reality (Li et al., 2022), robotics (Nasiriany et al., 2024) and many other applications. Recent advances in computer vision and graphics (Mildenhall et al., 2021; Kerbl et al., 2023) and deep learning (Dosovitskiy, 2020; Caron et al., 2021; Oquab et al., 2023), combined with the availability of large datasets of 3D objects (Deitke et al., 2023, 2024; Yu et al., 2023; Chang et al., 2015), have made it possible to learn neural networks that can generate 3D objects from text, single images or a sparse set of views, and to do so in an feed-forward manner, achieving significantly faster speeds than distillation-based methods (Poole et al., 2022; Qiu et al., 2024; Chen et al., 2023; Wang et al., 2023b).
A particularly successful family of 3D generators are the ones based on sparse-views reconstruction (Xu et al., 2024c; Siddiqui et al., 2024; Li et al., 2024b; Zhang et al., 2024c; Tang et al., 2024a; Xu et al., 2024b; Xie et al., 2024a; Wang et al., 2024c; Hunyuan3D, 2024). Compared to single-image reconstructors, multi-view reconstruction models generally produce better 3D assets. This advantage arises because the multi-view images implicitly capture the object geometry much better, substantially simplifying the reconstruction problem. However, to generate a 3D object from text or a single image, one must first synthesize several views of the objects, for example by means of a multi-view diffusion model. These multi-view diffusion models often generate inaccurate and inconsistent views, which are difficult for the reconstruction network to reconcile, and can thus affect the overall quality of the final 3D output (Tang et al., 2024c).
@@
In summary, this paper makes the following key contributions: (1) We introduce a novel pipeline that generates a pool of 2D views of an object and only selects the optimal subset for 3D reconstruction. (2) We propose FlexRM, a 3D reconstruction network that efficiently processes an arbitrary number of input views with varying viewpoints, enabling high-quality feed-forward reconstruction from the selected views. (3) We introduce a novel training strategy to enhance the robustness of the 3D reconstructor by simulating imperfect input views. This improves FlexRM’s resilience to small noise in the input data that may remain.
@@ 2.1 Multi-view Generation
Zero123 (Liu et al., 2023a) first proposed using multi-view data to fine-tune an image diffusion model for generating novel views from a single view, conditioned on camera parameters. Following this approach, subsequent works (Li et al., 2024b; Shi et al., 2023b; Tang et al., 2023; Liu et al., 2023b; Long et al., 2024; Wang & Shi, 2023; Woo et al., 2024; Yang et al., 2024b; Ye et al., 2024; Zhao et al., 2024; Tang et al., 2024b) largely focused on generating multiple views simultaneously to ensure 3D consistency.
@@ 2.2 Feed-forward 3D Reconstruction and Generation
Recent advances in 3D reconstruction and generation have focused on training feed-forward models that directly output 3D representations without requiring further optimization (Yu et al., 2021; Erkoç et al., 2023; Szymanowicz et al., 2023b; Ren et al., 2023; Lorraine et al., 2023; Xu et al., 2024a; Tochilkin et al., 2024; Zhang et al., 2024a; Jiang et al., 2024; Han et al., 2024a; Siddiqui et al., 2024). These feed-forward models offer significant advantages in both reconstruction quality and inference speed.
@@
A representative series of work is LRM (Hong et al., 2024), which learns to generate a tri-plane NeRF (Chan et al., 2022) representation using a transformer network. This approach can receive multiple types of inputs, including single images, text (Xu et al., 2024d), posed sparse-view images (Li et al., 2024b), and unposed sparse-view images (Wang et al., 2024a). Further works (Wei et al., 2024; Xu et al., 2024b; Wang et al., 2024c; Boss et al., 2024) focused on improving the geometric quality of generated 3D assets. Some (Zou et al., 2023) proposed to combine the tri-plane representation with 3D Gaussian Splatting for more efficient rendering. They suggest using an additional point cloud network to determine the 3D Gaussian position to overcome the tendency of 3D Gaussian to get stuck in local optima.
@@
Another representative series of work (Tang et al., 2024a; Xu et al., 2024c; Zhang et al., 2024c) generates 3D Gaussian points directly through per-pixel aligned Plücker ray embedding and predicts the depth for each pixel (Szymanowicz et al., 2023a), which can then be converted to 3D Gaussian locations. However, such approaches requires the input views to cover a large visible range of the 3D object. Building an intermediate 3D feature representation to regress 3D Gaussian points is also possible (Chen et al., 2024a; Zhang et al., 2024b), but these methods still require sparse-view images as input and typically prefer a fixed number of views with fixed viewing angles. However, in 3D generation tasks where sparse input views are generated through a multi-view diffusion model and are not always of high quality, such reconstructors tend to produce suboptimal results.
@@
This paper introduces FlexRM that combines the strengths of the approaches above. Our tri-plane-based model efficiently generates high-quality 3D Gaussian points directly, without needing additional modules. It also accommodates a variable number of input views.
2.3 Mitigating multi-view inconsistency with feedback.
Methods such as Ouroboros3D (Wen et al., 2024), Carve3D (Xie et al., 2024b), Cycle3D (Tang et al., 2024c), and IM-3D (Melas-Kyriazi et al., 2024) stem from similar motivations to our work, and they share a key idea: the feedback mechanism. While useful, these methods often require new supervision signals and learnable parameters to implement this feedback, potentially creating complex, monolithic pipelines that are difficult to decompose into reusable components for future designs. In contrast, Flex3D’s components are more easily generalized. Another key difference is Flex3D’s focus on the input to the reconstructor. This adds negligible computational cost and avoids the significant additional time required for multi-step refinement, preserving the speed advantage often associated with feed-forward models. Additionally, the feedback mechanism is orthogonal to our work and could be further combined.
@@ 3.1 Candidate View Generation and Selection
Our image/text-to-multi-view-images generator module generates four views of the 3D object from four elevation degrees (-18°, 6°, 18°, and 30°) We utilize the Emu model (Dai et al., 2023), which is pre-trained on a massive dataset of billions of text-image pairs, as our base model. Following prior works (Shi et al., 2023b; Li et al., 2024b; Siddiqui et al., 2024), we fine-tune this model on approximately 130,000 rendered multi-view images. This fine-tuning process enables the model to predict a 2$\times$2 grid of four consistent images, each corresponding to one elevation angle.
@@
As the two multi-view diffusion models are focused on different aspects (elevation and azimuth) of novel view generation, there are minimal conflicts in the generated views between them. Even so, and despite efforts to improve the quality of the outputs, not all generated views are entirely consistent. Certain views, particularly those from challenging angles like the back, may exhibit suboptimal generation quality, and there can be inconsistencies between different generated views. Including such flawed views as input for 3D reconstruction can significantly degrade the quality. Therefore, we introduce a mechanism to filtering poor-quality views. This is done via a novel view selection pipeline which consists of two steps:
(1) Back View Quality Assessment: We employ a multi-view video quality classifier trained to assess the overall quality of generated videos, with particular emphasis on the quality of the back view. This classifier utilized DINO (Oquab et al., 2023) to extract features from the front view and back view of the multi-view video, and subsequently trained a Support Vector Machine (SVM) to classify video quality based on the combined DINO features. The training data consisted of 2000 manually labeled “good” and “bad” Emu-generated video samples. We apply the quality classifier to the multi-view video to determine whether the back view exhibits reasonable generation quality.
@@
Candidate view generation typically takes about a minute on a single H100 GPU, and the whole process of view selection can be done in less than a second with a single A100 GPU.
@@ 3.2 Flexible Reconstruction Model (FlexRM)
The position of the 3D Gaussian is expressed as $\mathbf{p}=\alpha\mathbf{p}_{0}+(1-\alpha)\delta\mathbf{p}=\alpha\mathbf{p}_{0}+(1-\alpha)$ MLP$(f(\mathbf{p}_{0}))$ where $\delta\mathbf{p}$ is a a positional offset output by the MLP and $\alpha=3/4$. These positional offsets $\delta\mathbf{p}$ are constrained to the range of $[-1,1]$ through the application of the tanh activation function. This approach, akin to residual learning, facilitates the optimization process. The multipliers ensure that $\mathbf{p}$ remains within the same range as $\mathbf{p}_{0}$, which prevents Gaussian points from moving beyond the visible boundaries, which would result in their projection falling outside the 2D image plane and consequently providing no gradients for optimization. Furthermore, since $\alpha<1$, this expression biases Gaussians to shift towards the center of the tri-plane grid, where the object is usually located.
@@
Our training dataset comprises multi-view dense renders from an internal dataset analogous to Objaverse. Specifically, for each object, we render 512$\times$512 resolution images from 256 viewpoints, uniformly distributed across 16 azimuth and 16 elevation angles. This process yields approximately 700,000 rendered objects, with 140,000 classified as high-quality. Furthermore, we leverage the Emu-video synthetic dataset (Han et al., 2024a), which consists of 2.7 million synthetic multi-view videos. Each video comprises 16 frames capturing a 360-degree azimuth range.
@@
More details on implementation and training of FlexRM are presented in the Appendix B.
Figure 3: Imperfect Input View Simulation Examples. We simulate different kinds of imperfect input views by feeding FlexRM’s output back as input and manipulating 3D Gaussian parameters.
@@ 3.3 Imperfect input view simulation
While FlexRM’s outputs naturally contain small imperfections, we also introduce random perturbations to FlexRM’s generated 3D Gaussian points during step (1) to simulate a wider range of imperfect inputs and promote greater diversity. We sample a randomly sized small cube from the large tri-plane cube and add noise with varying intensities to the 3D Gaussian parameters, excluding rotation. For example, adding noise to positions can simulate part-level 3D inconsistencies, while adding noise to color parameters can simulate color distortions. Adding noise to opacity results in a speckled or streaky appearance, while adding noise to scale leads to a blurring effect. Figure 3 illustrates these effects. During training, each effect is randomly used with a probability of 0.2. This combines them for greater diversity. Please check Appendix B for more details.
@@ 4 Experiments
We leverage 404 deduplicated prompts from DreamFusion (Poole et al., 2022) to conduct an experiment on text-to-3D or single-image-to-3D generation. We compare Flex3D to a few recent feed-forward 3D generation methods including OpenLRM (He & Wang, 2023; Hong et al., 2024), VFusion3D (Han et al., 2024a), LGM (Tang et al., 2024a), InstantMesh (Xu et al., 2024b), and GRM (Xu et al., 2024c). We also compare Flex3D with two recent direct 3D generation (diffusion) methods: LN3Diff (Lan et al., 2024) and 3DTopia-XL (Chen et al., 2024b). For GRM, we utilize its provided Instant3D’s multi-view diffusion model to generate input multi-view images, and for InstantMesh, we employ the default Zero123++ (Shi et al., 2023a) for text-to-input multi-view image conversion. For direct 3D diffusion methods, we use their single-image-to-3D generation pipeline.
@@
To further evaluate the overall quality of the generated content, we conducted a user study. Participants were presented with pairs of 360° rendered videos—one generated by Flex3D and one by a baseline model—and asked to select their preferred video. We randomly selected 40 prompts for evaluation. The corresponding 40 pairs of generated videos were then independently evaluated by five users, with each user assessing all 40 pairs. For each pair, we collect five results, and the majority preference was recorded as a win rate for Flex3D. Results are also shown in table 1, where a significant number of votes goes to our method for its high quality, regardless of the baselines used for comparison. This shows that our method clearly generates better 3D assets.
@@
[…]
OpenLRM 1 15.83 0.821 0.209 0.602 - -
VFusion3D 1 19.10 0.827 0.158 0.759 - -
FlexRM 1 21.21 0.862 0.125 0.832 - -
InstantMesh 4 21.33 0.859 0.133 0.809 1.372 0.841
GRM 4 25.03 0.899 0.102 0.869 1.496 0.866
FlexRM 4 25.55 0.894 0.074 0.893 1.205 0.878
FlexRM 8 26.33 0.897 0.069 0.906 1.188 0.881
FlexRM 16 26.51 0.902 0.068 0.911 1.182 0.884
FlexRM 24 26.65 0.905 0.067 0.915 1.175 0.886
FlexRM 32 26.77 0.907 0.066 0.919 1.169 0.888
Table 2: Reconstruction Performance on the GSO Dataset. FlexRM consistently outperforms other baselines, where it achieves the best results across different input view settings. CD (Chamfer Distance) values are multiplied by $10^{-2}$, and NC denotes Normal Correctness. Increasing the number of input views for FlexRM leads to improved reconstruction quality.
In table 2 we report the performance of FlexRM on the GSO reconstruction task, using varying numbers of input views, namely 1, 4, 8, and 16. We compare our results to several baseline methods, including single-view reconstruction models (LRM (He & Wang, 2023), VFusion3D (Han et al., 2024a)) and sparse-view reconstruction models (InstantMesh (Xu et al., 2024b), GRM (Xu et al., 2024c)). For the single-view setting, we use the input view at 0° azimuth and 10° elevation as input. For the 4-view setting, we use views at 0°, 90°, 270°, and 360° azimuth degrees, all at 10° elevation. For other numbers of views, we heuristically select more views as input. The remaining views are used to compute the reported novel-view synthesis quality. The CD (Chamfer Distance) and NC (Normal Correctness) calculation protocol follows that of (Siddiqui et al., 2024).
Overall, FlexRM significantly outperforms baselines in both 1-view and 4-view settings. This improvement is particularly evident in the LPIPS score, a key metric reflecting perceptual quality, which demonstrates substantial gains. Beyond fixed input views, FlexRM is also capable of handling an arbitrary number of input views. We present results for more view results, both exhibiting progressively stronger performance. Qualitative results are shown in the Appendix D.